Calculate the resistance of copper wire with a diameter of 2.5968 mm having a length of 1000 m at 20°C.


1
Expert's answer
2021-05-17T17:01:01-0400


Resistivity of copper at 20"^0 C" is = 1.68 "\times 10^{-8} \Omega m"


Resistance = Resistivity "\times \frac{Length}{Area}"


R= "\rho \frac{L}{A}"


A= "\pi r^2" = 3.142 "\times (\frac{ 2.5968\times 10^{-3}}{2})= 5.2969\times 10^{-6} m^2"

R= 1.68 "\times 10^{-8} \times \frac{1000}{ 5.2969\times 10^{-6}}" = 3.1717 "\Omega"

R= "3.1717 \Omega"